OpenStudy (anonymous):

2x – y = 5 ? how do I do this?

OpenStudy (anonymous):

What are you trying to do?

OpenStudy (anonymous):

are you solving for y

OpenStudy (anonymous):

If you're trying to solve for y, then it is: 2x-y=5 subtract 2x -y=-2x+5 To make y positive, change all the signs to the opposite y=2x-5, and that's your answer.

OpenStudy (nincompoop):

this is an equation of a line there are usually 2 types of forms you can rewrite this into one is standard form, and the other is the y-intercept form standard form looks like this ax+by=c intercept form looks like this y=mx + b where m is your slope and b is your y-intercept so for example, you were given 5x - 2y = 5 you want to isolate y by itself, and preferably on the left by adding (-5x) on both sides of the equation. this becomes 5x + (-5x) - 2y = 5 + (-5x) -2y = -5x + 5 divide both sides by -2 y = 5x/2 - 5/2

OpenStudy (nincompoop):

If you want to convert 5x - 2y = 5 into standard form of a line. Now, if you recall the standard formula was ax + by = c and your 5x -2y = 5 is already in that format, so no conversion is necessary

OpenStudy (nincompoop):

@zeinas. please next time try not to give out answers even with worked out solutions try to provide a different problem and show steps solving it. giving answers to people will most likely make them skip the learning part and just copy the answer. this hinders the learning process and it is not efficient at all.

OpenStudy (anonymous):

thank you, everyone! I think I get it. quick question, if the answer to an equation like this was in a parentheses, would that mean I would have to solve it for both the y and the x? example: (2, –1)

OpenStudy (nincompoop):

no. knowing your slope, which is the m in the y-intercept form y = "m"x+b you can use the point-slope form y-y1 = m(x-x1) to solve for the actual equation of the line that passes through a specific point (x, y), and in your case (2, -1) x1 = 2 and y1 = -1

OpenStudy (anonymous):

Thanks!